Solve the system of equations by the addition method.
{ 5x -7y =4
{ 7x-3y=4













Select the correct answer


A) The solution is___?
B) There are infinitely many solutions
C) There is no solution

Answer by Alan3354(69443)

5x -7y =4
7x- 3y =4
------------
Multiply the 1st eqn by 3 and the 2nd by -7
---
15x - 21y = 12
-49x+ 21y = -28
-------------------------- Add
-34x = -16
================
solve for x, then sub and solve for y.
============
============
The addition method is multiplying 1 or both eqns by something to make the coefficients of one of the variables equal, then eliminating it.